OpenStudy (anonymous):

Question 7 (Multiple Choice Worth 1 points) Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). A. -9x - 2 B. -9x + 5 C.-9x + 25 D. -9x + 27

OpenStudy (jack1):

f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)) f(g(x)) = 9(-x + 3) - 2 f(g(x)) = -9x + 27 - 2 f(g(x)) = -9x + 25